# Human Data-Driven Calibration of Hand Exoskeletons Using Redundant Sensors for Improved Teleoperation

### v1

@@ III Human-Subject Experiments
Human-subject experiments were conducted to empirically optimize the calibration cost function weights and evaluate hand tracking performance. The study was approved by the Institutional Review Board (IRB ID: STUDY00002527). Seven participants with varying hand sizes (Table II) wore the exoskeleton and completed the protocol shown in Fig. 5. Reflective markers were placed on anatomical landmarks of the thumb and index finger for motion capture, as shown in Fig. 55.
@@
Second, during the hand tracking task, participants performed three representative finger motions: (i) combined thumb MCP and IP flexion, (ii) index MCP flexion with PIP extended, and (iii) combined index MCP and PIP flexion, as shown in Fig. 5. Each participant performed six trials per motion. These motions were selected to isolate individual joint estimation performance. Maestro sensor data and motion capture trajectories were recorded simultaneously and synchronized during post-processing. Due to limited visibility and occlusion when wearing the device, the analysis was limited to MCP and IP joints of the thumb, and MCP and PIP joints of the index finger. Due to the shared mechanical and kinematic design between the index and middle fingers, the results obtained for the index finger can also be applied to the middle finger.
@@ III-A Quantitative Analysis
To enhance calibration robustness against modeling uncertainties and local minima, cost function weights ($w_{k}$) were empirically optimized using a data-driven approach. Each subject performed dynamic finger movements while wearing the Maestro exoskeleton, and joint angle estimates were compared to ground-truth trajectories. For each subject, 500 candidate weight combinations ($w_{k}\in[0,10]$) were randomly sampled across all cost function dimensions. Each weight distribution was used to calibrate virtual link parameters from the two-phase calibration data, and compute the corresponding joint angles and fingertip positions. Motion capture trajectories served as ground truth to identify the weights that minimize the mean absolute error (MAE) in joint angle tracking. The weight distribution yielding the lowest error per subject was selected as optimal. Final weights were obtained by averaging these optimal weights across all participants and applied consistently in subsequent analyses.
@@
Figure 5: Experimental setup showing motion capture markers placed on the (a) thumb and (b) index finger, along with (c) the experiment workflow for human subject testing.
@@ III-B Qualitative Analysis
After verifying the quantitative performance of our exoskeleton-based hand tracking system, we qualitatively assessed how well the calibrated virtual hand resembled the user’s real hand motion in a virtual environment. A custom Unity hand model was developed using direct linear joint-to-joint mapping from the Maestro-estimated angles. DIP joints were inferred from corresponding PIP angles via biomechanical coupling, and all other joints were fixed. The visualization used anatomically plausible joint limits and averaged bone lengths across subjects and conditions, to isolate the effect of calibration. Participants reproduced representative static and dynamic hand gestures—–such as pinching and grasping–—under calibrated and uncalibrated conditions. These gestures were used to generate side-by-side visualization for offline inspection.
@@ IV-A Quantitative Results
The cost function weights used in the calibration routine is search by averaging the individual optimal weights distribution. The individual optimal weightsdistribution is from each subject by minimizing the joint angle mean absolute error relative to motion capture ground truth. The individual weight distributions for both the thumb and index fingers across all seven subjects, along with the averaged optimal weights, are shown in Fig. 6. While the individual profiles exhibit broadly consistent contours, indicating robustness in the optimal weights distribution, some outliers are observed. These deviations likely stem from subject-specific differences in hand size, donning variability, or movement variability during the calibration trials. The averaged results for the thumb (Fig.6) indicate higher optimal weights for $w_{2}$ and $w_{8}$, suggesting that the calibration is more sensitive to parameters contributing to $\theta_{2}$ and $\delta_{2}$, both related to the thumb MCP joint. In contrast, the index finger results (Fig.6) show greater weights for $w_{3}$ and $w_{4}$, highlighting the sensitivity of the calibration to $\delta_{1}$ and $\delta_{2}$, associated with the PIP joints of the index and middle fingers. Overall, the optimized weights emphasize the second kinematic loop parameters—particularly the thumb MCP and finger PIP joints—as key contributors to accurate hand tracking.
@@
We report mean absolute errors (MAE) between the Maestro-based tracking and ground truth for joint angles and fingertip positions. Average results are shown in Fig. 7, while subject-specific results for the three tracking tasks are shown in Fig. 8. Optimal-weighted calibration consistently achieved the lowest MAE, followed by the even-weighted calibration, with the uncalibrated model exhibiting the highest error. This trend holds across joints for the tested digits, validating the benefit of subject-specific weight optimization.
Subject Length Width Size Joint MAE $\downarrow$ [%] Tip MAE $\downarrow$ [%]
[cm] [cm] T-MCP T-IP I-MCP I-PIP Thumb Index
1 17.0 6.5 S 8.5-8.5- 8.5 64.764.764.7 53.553.553.5 64.364.364.3 32.232.232.2 69.069.069.0
2 17.2 8.0 S–M 39.5-39.5- 39.5 11.411.411.4 70.870.870.8 44.244.244.2 5.35.35.3 60.760.760.7
3 18.0 9.5 M–L 6.6-6.6- 6.6 12.512.512.5 40.5-40.5- 40.5 57.057.057.0 32.132.132.1 58.058.058.0
4 18.5 9.0 M 33.633.633.6 48.648.648.6 3.1-3.1- 3.1 61.761.761.7 25.225.225.2 63.863.863.8
5 17.0 8.5 M 18.4-18.4- 18.4 3.23.23.2 64.264.264.2 43.943.943.9 41.241.241.2 85.285.285.2
6 19.2 9.8 L 59.159.159.1 66.166.166.1 73.673.673.6 56.456.456.4 72.172.172.1 83.383.383.3
7 18.1 7.8 M 10.8-10.8- 10.8 62.162.162.1 57.257.257.2 35.435.435.4 3.03.03.0 66.066.066.0
Avg 17.6 8.4 37.1 56.7 68.3 52.9 34.8 71.5
TABLE II: Hand dimensions (Length, Width in cm) and percent reduction in joint and fingertip tracking errors after calibration.
@@
Overall, the index finger showed the most consistent improvements across subjects, particularly in PIP joint tracking and fingertip accuracy. The thumb also benefited from calibration, though with greater variability—likely due to anatomical differences and higher articulation complexity. A few negative joint error reductions (i.e., increases in error post-calibration) were observed in the thumb MCP and index MCP joints for certain subjects, possibly due to residual alignment errors during marker placement or mechanical coupling not fully captured by the model. Despite inter-subject variability, the averaged error reductions and similarity in weight profiles suggest that the calibration method generalizes well across users with different hand sizes.
Figure 7: Hand tracking accuracy results averaged across 7 subjects and all tasks, shown for (left) joint angles (thumb MCP, thumb IP, index MCP, index PIP) and (right) fingertip position (thumb, index).
@@
Figure 8: Hand tracking accuracy results for 7 subjects. (a) Thumb joint angles and fingertip errors across all subjects. (b) Index joint angles and fingertip errors across all subjects for two different index finger motions.
To illustrate the method’s performance, Fig. 9 shows tracking results for Subject 6 across thumb and index. The weighted calibration consistently produces trajectories that align more closely with the ground truth data, compared to both the uncalibrated and even-weighted baselines. This subject serves as a representative example of the effectiveness of our approach in reducing tracking error in dynamic finger motion.
@@
Figure 9: Comparison of hand tracking performance for Subject 6. Each subplot compares trajectories obtained from motion capture, uncalibrated tracking, and calibration with even and optimal weights.
@@ IV-B Qualitative Results
To qualitatively assess improvements in hand tracking, we visualized hand tracking results using a virtual Unity hand driven by joint estimates under both calibrated and uncalibrated conditions. Representative postures (e.g., pinch, grasp) were recorded for each subject and used to qualitatively assess tracking fidelity. As shown in Fig. 10, orange hands depict postures based on uncalibrated data, while green hands reflect calibrated estimates.
Calibration leads to visibly improved alignment with the real hand, particularly in fine thumb and finger articulations. These visualizations qualitatively confirm the enhanced motion fidelity achieved through the proposed subject-specific calibration. A demonstration of the real-time virtual hand motion tracking is provided in the supplementary video.
@@
V Discussion and Conclusion
This study introduced a subject-specific calibration framework to enhance hand tracking accuracy with a sensorized robotic exoskeleton. The method yielded consistent improvements in both joint angle and fingertip tracking across participants with varying hand sizes and donning conditions. We also introduced an optimal weight distribution for the calibration process, which consistently outperformed even-weighted calibration, particularly in joints affected by closed-loop kinematics. These results suggest that inaccuracies in virtual link parameters, originated from user-specific fit, can significantly affect tracking and should be explicitly calibrated. In a few cases, such as Subject 3, improvements were not observed, likely due to external factors like motion capture occlusion or soft tissue artifacts. Moreover, when initial tracking errors were already low, optimization converged near a local minimum, yielding marginal or oscillatory improvements.
When compared to existing tracking systems, our method demonstrated competitive performance. Markerless vision-based methods often report fingertip errors exceeding 10–20mm[23], while our calibrated system achieves fingertip MAE of 10mm—sufficient for teleoperation applications [12]. Compared to other exoskeleton-based approaches that rely on structured calibration routines [24], our method avoids constrained motion patterns and demonstrates more robust tracking across users. A key contribution is the validation of performance across a diverse user group, in contrast with prior single-subject evaluations. We also provide both quantitative and qualitative evidence of improved motion plausibility using a Unity-rendered virtual hand.
Importantly, the proposed calibration framework is generalizable. Many hand exoskeletons employ 4-bar linkages or underactuated designs due to the challenge of aligning robotic joints with anatomical joints [15]. These designs require virtual kinematic models whose geometry is affected by unmeasurable donning conditions. Our approach compensates for such inaccuracies by optimizing virtual link lengths to minimize joint estimation error. While additional sensors improve observability and constraint in the optimization, our method is compatible with systems lacking redundancy. By addressing anatomical variability and user-specific fit, this work helps bridge the gap between human intent and virtual or robotic hand motion.
Despite the promising results, this study presents several limitations. First, the evaluation was limited to isolated postures rather than continuous movements, primarily due to motion capture occlusion issues. While this enabled precise joint-level analysis, it did not fully capture the temporal complexity of real-world manipulation tasks. Second, the calibration assumes a consistent exoskeleton fit and stable sensor behavior, yet in practice, variations in donning, long-term use, and environmental conditions may affect performance and reduce calibration validity over time.
In conclusion, this work lays the foundation for extending subject-specific calibration methods to practical teleoperation and robot learning applications. By combining subject-specific virtual links optimization with a quantitative validation, we present a framework to enhance the fidelity of exoskeleton-based hand tracking. Qualitative assessments using a simulated teleoperation hand model confirmed that calibration leads to more anatomically plausible postures, and showed improved visual alignment post-calibration. Although residual joint angle errors remain in the 10-degree range, the resulting motions are visually coherent and consistent. Such improvements in motion fidelity are essential for downstream teleoperation tasks involving fine manipulation, shared autonomy, or learning from demonstration. Future work will focus on evaluating performance in dynamic, task-based scenarios and integrating the framework into imitation learning pipelines to assess its effect on policy generalization and data efficiency.

### v2

@@ III Human-Subject Experiments
Human-subject experiments were conducted to empirically optimize the calibration cost function weights and evaluate hand tracking performance. The study was approved by the University of Texas Institutional Review Board (IRB ID: STUDY00002527). Seven participants with varying hand sizes (Table II) wore the exoskeleton and completed the protocol shown in Fig. 5. Reflective markers were placed on anatomical landmarks of the thumb and index finger for motion capture, as shown in Fig. 5-5.
@@
Second, during the hand tracking task, participants performed three representative finger motions: (i) combined thumb MCP and IP flexion, (ii) index MCP flexion with PIP extended, and (iii) combined index MCP and PIP flexion, as shown in Fig. 5. These motions were selected to isolate individual joint estimation performance and were performed six times. Maestro sensor data and motion capture trajectories were recorded simultaneously and synchronized during post-processing. Due to limited visibility and occlusion when wearing the device, the analysis was limited to MCP and IP joints of the thumb, and MCP and PIP joints of the index finger. Because the index and middle fingers share an identical mechanical and kinematic design, calibration and analysis were performed on the index finger only.
@@ III-A Quantitative Analysis
To enhance calibration robustness against modeling uncertainties and local minima, cost function weights ($w_{k}$) were empirically optimized using a data-driven approach. Each subject performed dynamic finger movements while wearing the Maestro exoskeleton, during which a random search sampled candidate weight combinations for the cost function, and the resulting joint angle estimates were compared against ground-truth trajectories. For each subject, 500 candidate weight combinations ($w_{k}\in[0,10]$) were randomly sampled. Each weight distribution was used to calibrate virtual link parameters from the two-phase calibration data, and compute the corresponding joint angles and fingertip positions. Motion capture trajectories served as ground truth to identify the weights that minimize the mean absolute error (MAE) in joint angle tracking. The weight distribution yielding the lowest error per subject was selected as optimal. Final weights were obtained by averaging these optimal weights across all participants and applied consistently in subsequent analyses.
@@
Figure 5: Experimental setup and workflow. (a) Motion capture markers placement on the thumb, (b) on the index finger, and (c) experiment workflow for human subject testing.
@@ III-B Qualitative Analysis
After verifying the quantitative performance of our exoskeleton-based hand tracking system, we qualitatively assessed how well the calibrated virtual hand resembled the user’s real hand motion in a virtual environment. A custom Unity hand model was developed using direct linear joint-to-joint mapping from the Maestro-estimated angles. DIP joints were inferred from corresponding PIP angles via biomechanical coupling, and all other joints were fixed. The visualization used anatomically plausible joint limits and averaged bone lengths across subjects and conditions, to isolate the effect of calibration. Participants reproduced representative static and dynamic hand gestures (e.g., pinching) under calibrated and uncalibrated conditions. These gestures were used to generate side-by-side visualization for offline inspection.
@@ IV-A Quantitative Results
The cost function weights used in the calibration routine were obtained by averaging subject-specific optimal distributions, each derived by minimizing the mean absolute joint angle error relative to motion capture ground truth. The resulting thumb and index finger distributions across all seven subjects, along with the averaged weights, are shown in Fig. 6. While the individual profiles exhibit broadly consistent contours, indicating robustness in the optimal weights distribution, some outliers are observed. These deviations likely stem from subject-specific differences in hand size, donning variability, or movement variability during the calibration trials. The averaged results for the thumb (Fig. 6) indicate higher optimal weights for $w_{2}$ and $w_{8}$, suggesting that the calibration is more sensitive to parameters contributing to $\theta_{2}$ and $\delta_{2}$, both related to the thumb MCP joint. In contrast, the index finger results (Fig. 6) show greater weights for $w_{3}$ and $w_{4}$, highlighting the sensitivity of the calibration to $\delta_{1}$ and $\delta_{2}$, associated with the PIP joints of the index and middle fingers. Overall, the optimized weights emphasize the second kinematic loop parameters—particularly the thumb MCP and finger PIP joints—as key contributors to accurate hand tracking.
@@
We report mean absolute errors (MAE) between the Maestro-based tracking and ground truth for joint angles and fingertip positions. Average results are shown in Fig. 7, while subject-specific results for the three tracking tasks are shown in Fig. 8.
Sub Length Width Thumb (MAE %) Index (MAE %)
MCP IP Tip MCP PIP Tip
(cm) (cm) Even Opt. Even Opt. Even Opt. Even Opt. Even Opt. Even Opt.
1 17.0 6.5 -74.7 -9.3 50.7 64.6 -16.3 8.7 59.1 57.2 33.2 54.3 60.7 78.1
2 17.2 8.0 2.2 -25.5 11.4 11.4 14.7 5.3 77.7 77.9 66.4 70.0 11.2 22.3
3 18.0 9.5 2.3 -6.6 37.5 12.5 23.3 32.1 -71.3 -39.0 63.6 75.6 -34.4 19.3
4 18.5 9.0 -48.3 43.6 48.6 48.6 -87.8 25.2 -0.8 -3.1 35.4 56.1 81.1 75.6
5 17.0 8.5 -29.4 -15.4 -2.9 3.2 23.6 41.2 39.3 42.9 45.0 47.3 78.5 85.2
6 19.2 9.8 45.8 69.1 12.7 66.1 61.1 72.1 70.0 74.5 61.0 69.5 70.6 82.4
7 18.1 7.8 -7.9 -10.8 62.1 62.1 5.3 3.0 58.2 58.2 39.0 53.6 77.1 78.3
Avg 17.6 8.4 26.9 37.1 40.4 56.7 44.2 68.3 42.7 52.9 12.7 34.8 58.6 71.5
TABLE II: Hand dimensions and calibration results for seven subjects. Entries show the percent change in joint/tip MAE under even-weighted (Even) and optimal-weighted (Opt.) calibration relative to the uncalibrated baseline.
@@
The optimal-weighted calibration consistently yielded the lowest MAE, followed by the even-weighted calibration, while the uncalibrated model exhibited the highest errors across all tested joints. On average, the optimal-weighted approach produced greater error reductions, particularly at the fingertip level, with the index finger showing the most consistent improvements. The thumb also showed improvement following calibration, albeit with greater variability, which can be attributed to anatomical variability and the complexity of its joint articulation. Occasional negative error reductions in the thumb and index MCP joints may reflect residual alignment errors or unmodeled mechanical coupling. Overall, the consistent error reductions and comparable optimal weight profiles across participants indicate that the proposed calibration method generalizes well to users with varying hand sizes.
Figure 7: Hand tracking accuracy results averaged across subjects and tasks. (left) thumb and index joint angles errors and (right) thumb and index fingertip position errors.
@@
Figure 8: Hand-tracking accuracy results for 7 subjects. (a) Thumb MCP and IP joint angle errors and fingertip position errors (b) Index-finger MCP and PIP joint angle errors and fingertip position errors.
As shown in Fig. 9, for Subject 6, the weighted calibration achieves trajectories that align most closely with the ground truth, outperforming both the uncalibrated and even-weighted models. This example highlights the method’s effectiveness in reducing tracking errors during dynamic finger motion.
@@
Figure 9: Comparison of hand-tracking performance for Subject 6. Motion-capture trajectories are shown alongside uncalibrated and calibrated estimates using even and optimal weights.
@@ IV-B Qualitative Results
To qualitatively assess tracking improvements, we visualized results using a virtual Unity hand, driven by joint estimates under calibrated and uncalibrated conditions. Representative postures (e.g., pinch) were recorded for each subject to visually evaluate tracking fidelity. As shown in Fig. 10, red hands represent uncalibrated data, while green hands show calibrated poses. Calibration visibly improves alignment with the real hand, especially in fine thumb and finger movements, qualitatively confirming the enhanced motion fidelity. A demonstration of real-time virtual hand tracking is provided in the supplementary video.
@@
V Discussion
This study introduced a subject-specific calibration framework to improve hand-tracking accuracy in a sensorized robotic exoskeleton. As shown in Section IV, calibration reduced both joint angle and fingertip errors across participants with varying hand sizes and donning conditions, confirming the method’s effectiveness. An optimal weight distribution further improved performance, particularly in joints influenced by closed-loop kinematics, confirming that user-specific variability and misalignment must be explicitly calibrated by prioritizing kinematic parameters that are more sensitive to different hand sizes or slippage.
Previous studies have also examined calibration trade-offs and shape-specific adaptation in data-glove systems [19, 20], which align with our findings and further emphasize the need for efficient, subject-agnostic calibration strategies.
Our framework applies this concept to robotic exoskeletons by embedding subject-specific calibration directly into the device’s kinematic model. By optimizing virtual parameters from redundant sensing and reference hand poses, it achieves subject-specific kinematic alignment without additional instrumentation or repetitive manual calibration.
In terms of performance, the proposed method achieves a mean absolute fingertip error of approximately 10 mm, comparable to or better than markerless vision-based systems that often exceed 10–20 mm [25]. Prior studies have reported tolerable fingertip errors up to 10-15 mm without perceptible loss of control fidelity [26, 12], supporting the feasibility of the proposed framework for teleoperation tasks.
While the calibration approach is demonstrated on the Maestro hand exoskeleton, the proposed calibration framework extends naturally to other wearable devices where tracking depends on human parameters and there are sufficient independent constraints. For instance, in this study, the optimization was performed on kinematic loops lacking redundant sensing, such as the thumb IP joint, where closed-loop geometric constraints were incorporated to provide sufficient constraints. The data-driven weight optimization generalizes across users, as similar weight distributions were observed among participants, allowing an averaged profile to be used. Motion-capture calibration may only be needed once per device to account for design-specific geometry and sensors. The only fundamental requirement is that the number of virtual link parameters to be optimized does not exceed the number of independent kinematic constraints derived from redundant sensing, geometric closure, or constrained reference poses.
Despite the promising results, this study presents several limitations. First, although the evaluation included continuous finger motion, the analysis primarily focused on short, isolated movement segments rather than long-duration or coordinated hand actions. While this enabled precise joint-level analysis, it did not fully capture the temporal complexity of real-world manipulation tasks. Second, we acknowledge that the participants represent hand sizes slightly below the population mean, highlighting the need for validation across a broader and more diverse range of users. Additionally, this evaluation focuses on relative error reduction rather than direct comparisons with other devices. Future work will include systematic benchmarking to establish unified accuracy standards for reliable hand-exoskeleton deployment in teleoperation. Future work will also explore online or self-adaptive calibration schemes, allowing real-time adjustment to changes in fit or sensor drift. Extending the framework to simultaneous multi-finger calibration could further enhance dexterity modeling and real-world usability across diverse robotic hands.
VI Conclusion
In conclusion, this work presents a subject-specific calibration framework that improves the accuracy of exoskeleton-based hand tracking for teleoperation. Implemented on the Maestro hand exoskeleton, the framework significantly reduces tracking errors by embedding anatomical variability derived from representative users directly into the device’s kinematic model. By optimizing virtual link parameters and residual weights, it achieves consistently higher hand tracking fidelity. Although residual joint-angle errors of about 10° remain, the calibration enables precise tracking for fine manipulation and shared-autonomy tasks. The proposed method is readily generalizable to other hand-exoskeleton designs, supporting user-specific hand motion tracking.